OpenStudy (anonymous):

Use implicit differentiation to find dy/dx: ln xy + 5x = 30

OpenStudy (anonymous):

d/dx (lnxy +5x) = d/dx (30) d/dx(lnxy) + d/dx(5x)=0 d/dxy (lnxy) * d/dx (xy) + 5=0 1/xy ( y+ x dy/dx)=-5

OpenStudy (anonymous):

DOES IT HELP?

OpenStudy (anonymous):

Yes, haha, Thank you.